Planning Page for Feb D/Eng/Com Nights



Night List and Splits

  • UT Feb 2- PETS/DEC Split
    • DEC twilight-20:14hr (local)
    • PETS 1 - 20:15-02:50hr (local)
    • DEC 02:51hr-twilight (local)
    • See wiki page for plan - night lost to UPS issues and weather
  • UT Feb 3 - PETS/DEC Split
    • DEC twilight-22:29 (local)
      • LUCI-AO Testing
      • If LUCI-AO cannot be executed - LBC/TMS
    • PETS 2 - 23:00-03:30hr (local)
    • DEC 03:31-twlight (local)
      • LUCI-AO testing
    • See wiki page for details - night lost to UPS issues and weather
  • UT Feb 4 - PETS/DEC Split
    • DEC twilight-20:00 (local)
      • LUCI AO-testing
    • PETS 3 - 20:01-02:54hr (local)
    • DEC 02:55-twilight (local)
      • Reconfig PEPSI to LUCI - 02:55-03:07
      • Pointing/Collimation 03:08-03:23
      • LUCI-Bino SL - GRB 211211A - 03:24-04:39
        • Can be moved earlier if good conditions and AO not available (technical reasons)
    • See wiki page for details - night lost to UPS issues - testing UPS overnight
  • UT Feb 7 - PETS/DEC Split
    • DEC twilight-23:04hr (local)
      • LBTI - CQ Tau w/ALES+vAPP coronagraph sunset(17:57)-22:57 - 5hr after sunset
    • PETS 4- 23:25-04:16hr (local)
    • DEC 04:17hr-twilight (local)
    • Poor seeing whole night - only TMS testing acquired any data
  • UT Feb 8 - PETS/DEC Split
    • PETS-5 - twilight-00:45hr (local)
    • DEC 00:46hr-twilight (local)
    • Fairly successful LUCI-AO testing
  • UT Feb 14 - Full DEC night
    • LBC/TMS - 4hrs - twilight -23:30
    • LUCI-1 AO/ESM - 00hr-twilight
    • Successful LBT/TMS testing - lots of AO/ESM testing
  • UT Feb 15 - Split w/LBTI
    • LBTI 1st half
    • Night Center 00:33 (local)
    • DEC 2nd half
      • LBTI LMIRCam photometric standards 3hr 00:33-03:33
      • Reconfigure from LBTI to bino-MODS 03:34-03:49
      • MODS Engineering tasks
        • run zernike_collect to measure rotator offset angle for active optics -1.5hr
        • if clear - take a full set of settings for a spectro-photometric standard
    • Tasks completed
  • UT Feb 16 - Split w/LBTI
    • LBTI 1st half
    • Night Center 00:33 (local)
    • DEC 2nd half
    • Two brief chances to open in between gusts - couldn't even complete collimation
  • UT Feb 18 - Split w/LBTI
    • LBTI 1st half
    • Night Center 00:33 (local)
    • DEC 2nd half
    • Highly variable seeing - some AO attempted - Moon angular separation tests, Bin 2 ESM does not rollback to Bin 1, rolled edges seen Bin 1 Full AO in unstable conditions, tested AO supervisor update (unsure if it works...TBD)
  • UT Feb 20 - PETS 6 05:50-10:10 UT (22:50-03:10 local)
    • DEC twilight-22:20 (local)
    • DEC 03:11-twilight (local)
  • Handover/Reconfigure time:
    • PETS 30min before the science critical start time
    • LBTI - split the reconfigure time between DEC and LBTI - exact time depends on instrument used by DEC

Engineering Tasks

  • LUCI-1 AO/ESM Testing
    • on-axis full AO Bin 1 and Bin 3
    • off-axis or on-axis "pretty pix" - any Bins
    • ESM Bin 1 and Bin 2
    • quantify the impact of moonlight - test acquisition at varying angular distances from the Moon
  • LBC TMS testing (3-4 hrs per block)
  • DONE LBTI LMIRCam photometric standards (3hr, seeing <1.5 arcsec, thin clouds at most, after UT Feb. 14 for UBC cold)
  • Photometric Standards for LBCs, LUCIs, MODS (1-3 hr - poor seeing and photometric conditions only)
    • LUCI - /home/lbto/Service_Observing/2022A_D_E/PoorWeather/LUCI_Phot_Standards/LUCI_Phot_ZP_OT_V3.xml
    • If seeing > 4" - make these track modes in the OT and export
  • DONE MODS - run zernike_collect to measure rotator offset angle for active optics -1.5hr

Science Programs

  • LBTI - CQ Tau w/ALES+vAPP coronagraph - 5hr (minimum: 4h after sunset, beginning of the night, clear, seeing <1.2 arcsec) - Ertel
  • LUCI-1 AO - TIC 266958963 R=14.6 Bin 3 - 0.5hr (1st half of night) - Zhang
    • pending release of Bin 3 Full AO - will be scheduled
  • LUCI-Bino SL - GRB 211211A - 1.25hr (end of night) - Rothberg
  • LBC - Comet Halley (mag 25.5) - 1.5hrs
    • 2x45 min periods - few hours apart (to demonstrate motion)
    • Seeing <1.2" thin clouds - dark
    • g and r filters
    • as of 2/4/22 RA 08:18 Dec+02:11 mag=25.5
    • For NSIGUI use "90000033"


-- BarryRothberg - 27 Jan 2022
Topic revision: r22 - 18 Feb 2022, BarryRothberg
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback